OpenStudy (anonymous):

plzzzzz help me. . . . The half-life of a radioactive substance is 5 minutes. how much of an original mass of 100 grams remains after half an hour?

OpenStudy (anonymous):

but the choices are a=14825 b. 24685 c. 15625 d. 25625

OpenStudy (anonymous):

ok let me try after first half lif you will have 50 grams after 30 minutes (6 half lives ) =100*(0.5)^6=1.5625

OpenStudy (anonymous):

if \[\log_{36.2} =1.5587\]

OpenStudy (anonymous):

what is the characteristics of\[\log=36.2\]